Discrete Mathematics.

Question 2:
(a) Use modular arithmetic to find 1040 mod 210. Show your working.
(b) An RSA cryptosystem uses public key pq = 65 and e = 7. Decrypt the ciphertext 57 9 and translate the result into letters of the alphabet to discover the message.
Answer #1

we have to find loto mod 210 105 10 (mod 21. ) 10²= 100 (mod 210 ) 103 1000 = 160 (mod 2 10 ) 1600 130 (mod 210) > (mod 2 loput K=1 » da 1x (6-1)(13-1) + 4.*12+1 49 7 7 → 10=7 Now MIE 8 (57) (mod 65) M2 = 91? (mod65) - 9 & > Mi=8 - H 1- In Englishif you have any doubt then feel free to ask me.

Please like my solution.